OpenStudy (anonymous):

What are the x-intercept(s) of the graph of y + 5x = x2 − 24? (12, 0) and (−2, 0) (−12, 0) and (2, 0) (8, 0) and (−3, 0) (−8, 0) and (3, 0)

OpenStudy (johnweldon1993):

\[\large y + 5x = x^2 - 24\] We want this to look like \[\large y = ax^2 + bx + c \] So lets subtract 5x from both sides of the equation \[\large y = x^2 - 5x - 24 \] Now we can use the quadratic formula or complete the square to find the zeros (x-intercepts) of the equation

OpenStudy (anonymous):

(x + 3)(x - 8)

OpenStudy (triciaal):

so x = -3 x = 8 points (-3, 0) and (8,0) same as choice C

OpenStudy (johnweldon1993):

Perfect :)